# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 3.3.1. The Donaldson’s functional
(3.38) $\|s\|_{P}=\sup_{X\setminus\Sigma}|s|+\|\bar{\partial}_{\theta}s\|_{L^{2}}+\|\Delta_{\omega,K}^{\prime}s\|_{L^{1}}.$
Let $\mathcal{P}$ denote the space of smooth Hermitian metrics $K$ such that $\int_{X\setminus\Sigma}|\Lambda_{\omega}F_{K,\theta}|_{\omega}\frac{\omega^{n}}{n!}<\infty.$ Due to Proposition 4.1 in [97], we can make $\mathcal{P}$ into analytic manifold with local charts $s\mapsto Ke^{s}$, $s\in\mathcal{P}(S_{K})$. Let $\mathcal{P}_{0}$ be one of the components covered by these charts. Let’s recall the Donaldson’s functional defined on the space $\mathcal{P}_{0}$ (see Section 5 in [97] for details),
@@
The following proposition is standard because the argument of [97, Section 5] can be directly applied since $d\omega^{n-1}=0$.
@@
Let $H(t)=Kh(t)$, $0\leq t\leq 1$, be a piecewise differentiable curve connecting metrics $K$ and $H$ in the space $\mathcal{P}_{0}$. If $d\omega^{n-1}\equiv 0$, then
(3.40) $\mu_{\omega}(K,H)=\int_{0}^{1}\int_{M\setminus\Sigma}\sqrt{-1}\operatorname{tr}(h^{-1}\frac{\partial h}{\partial t}\cdot F_{H(t),\theta})\wedge\frac{\omega^{n-1}}{(n-1)!}dt,$
@@
(3.41) $\frac{d}{dt}\mu_{\omega}(K,H(t))=\int_{M\setminus\Sigma}\sqrt{-1}\operatorname{tr}(h^{-1}\frac{\partial h}{\partial t}\cdot F_{H(t),\theta})\wedge\frac{\omega^{n-1}}{(n-1)!}$
@@ 3.3.2. Approximate Hermitian-Einstein structure
Let $H(t)$ be the long time solution of (3.9) on the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})|_{X\setminus\Sigma}$. Since we have known that $|\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}|_{H(t)}$ is uniformly bounded for $t\geq t_{0}>0$, by (3.37) and Lemma 3.17, it is easy to see that $H(t)$ (for every $t>0$) belongs to the space $\mathcal{P}_{0}$. By (3.41), we have a formula for the derivative of Donaldson’s functional with respect to $t$,
@@
(3.43) $r^{-\frac{1}{2}}\|s(t_{1},t_{2})\|_{L^{1}(X\setminus\Sigma,\omega,H(t_{1}))}\leq 2\hat{C}_{1}(t_{2}-t_{1})+\operatorname{Vol}(M,\omega)\log(2r),$
(3.44) $\|s(t_{1},t_{2})\|_{L^{\infty}(X\setminus\Sigma,H(t_{1}))}\leq r^{\frac{1}{2}}\{2C_{S}\hat{C}_{1}(t_{0}^{-n}+1)(t_{2}-t_{1})+\log 2r\}$
@@
(3.45) $\|s(t_{1},t_{2})\|_{L^{\infty}(X\setminus\Sigma,H(t_{1}))}\leq C_{1}(t_{0}^{-1})\|s(t_{1},t_{2})\|_{L^{1}(M\setminus\Sigma,\omega,H(t_{1}))}+C_{2}(t_{0}^{-1})$
for $0<t_{0}\leq t_{1}\leq t_{2}$, where $C_{1}(t_{0}^{-1})$ and $C_{2}(t_{0}^{-1})$ are constants depending only on $r$, the Sobolev constant $C_{S}$, $\hat{C}_{1}$, $t_{0}^{-1}$ and $\operatorname{Vol}(X,\omega)$.
@@
(3.46) $\int_{X\setminus\Sigma}|\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta_{X_{\mathrm{reg}}}}-\lambda\operatorname{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}|_{H(t)}\frac{\omega^{n}}{n!}\leq\hat{C}_{1},$
where $\hat{C}_{1}$ is a positive constant. Setting
(3.47) $\int_{X\setminus\Sigma}|\sqrt{-1}\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}-\lambda\operatorname{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}|_{H(t)}^{2}\frac{\omega^{n}}{n!}\rightarrow C^{\ast}\geq 0,$
as $t\rightarrow+\infty$. Assume that there is a sequence $t_{i}$ such that
@@
Then the following statement holds.
  1. (i)
    There is a section $)u_{\infty}\in L_{1}^{2}(\Gamma(S_{\hat{H}}$)) such that $\operatorname{tr}u_{\infty}=0$, $\|u_{\infty}\|_{L^{2}}=1$ and the eigenvalues of $u_{\infty}$ are almost everywhere constants and not all equal. Furthermore, assume $\mu_{1}<\mu_{2}<\cdots<\mu_{l}$ ($l\geq 2$) are the distinct eigenvalues of $u_{\infty}$. Defining smooth functions $P_{\alpha}:\mathbb{R}\to\mathbb{R}$ such that
    (3.49) $P_{\alpha}(x)=\left\{\begin{split}\ 1,\quad&x\leq\mu_{\alpha},\\ \ 0,\quad&x\geq\mu_{{\alpha}+1}.\end{split}\right$.
    and setting $\pi_{\alpha}=P_{\alpha}(u_{\infty})$ for every $1\leq\alpha\leq l-1$. Then every $\pi_{\alpha}$ defines a saturated $\theta_{X_{\mathrm{reg}}}$-invariant subsheaf $\mathcal{E}_{X_{\mathrm{reg}},\alpha}$ of the Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$, and every $\mathcal{E}_{X_{\mathrm{reg}},\alpha}$ can be extended to a saturated subsheaf $\mathcal{E}_{\alpha}$ of $\mathcal{E}_{X}$.
  2. (ii)
    $2\pi\sum_{\alpha=1}^{l-1}(\mu_{\alpha+1}-\mu_{\alpha})\operatorname{rank}(\mathcal{E}_{\alpha})(\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})-\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{\alpha}))\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.$
@@
(3.50) $\lim_{i\rightarrow\infty}\|s(t_{0},t_{i})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}\rightarrow+\infty,$
@@
for all $0<t_{0}\leq 1\leq t$, where $C_{3}$ and $C_{4}$ are constants depending only on the Sobolev constant $C_{S}$, $r$, $\hat{C}_{1}$ and the geometry of $(M\setminus\Sigma,\omega)$.
Setting $u_{i}(t_{0})=\|s(t_{0},t_{i})\|_{L^{1}}^{-1}s(t_{0},t_{i})\in S_{H(t_{0})}$, then $\|u_{i}(t_{0})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}=1$ and
@@
(3.54) $\liminf_{t\rightarrow+\infty}\frac{-\mu_{\omega}(H(t_{0}),H(t))}{\|s(t_{0},t)\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}}\geq r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.$
By the inequalities (3.54), (3.48), (3.51) and the Lemma 5.4 in [97], we can see that, by choosing a subsequence which we also denote by $u_{i}(t_{0})$, we have $u_{i}(t_{0})\rightarrow u_{\infty}(t_{0})$ weakly in $L_{1}^{2}$ as $i\rightarrow\infty$, where the limit $u_{\infty}(t_{0})$ satisfies: $\|u_{\infty}(t_{0})\|_{L^{1}(M\setminus\Sigma,\omega,H(t_{0}))}=1$, $\int_{M}\operatorname{tr}(u_{\infty}(t_{0}))\frac{\omega^{n}}{n!}=0$ and
@@
Furthermore, if $\Upsilon:R\times R\rightarrow R$ is a positive smooth function such that $\Upsilon(\lambda_{1},\lambda_{2})<(\lambda_{1}-\lambda_{2})^{-1}$ whenever $\lambda_{1}>\lambda_{2}$, then
(3.56) $\begin{split}&\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}(t_{0})\sqrt{-1}\Lambda_{\omega}(F_{H(t_{0}),\theta}))+\langle\Upsilon(u_{\infty}(t_{0}))(\overline{\partial}_{\theta}u_{\infty}(t_{0})),\overline{\partial}_{\theta}u_{\infty}(t_{0})\rangle_{H(t_{0})}\frac{\omega^{n}}{n!}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.\end{split}$
By (3.55) and (3.14), we known that $\|\Lambda_{\omega}(F_{H(t_{0}),\theta})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}$ and $\|u_{\infty}(t_{0})\|_{L^{\infty}(X\setminus\Sigma,H(t_{0}))}$ are uniformly bounded (independent of $t_{0}$). Then, (3.56) implies there exists a uniform constant $\check{C}$ independent of $t_{0}$ such that $\int_{X\setminus\Sigma}|\overline{\partial}_{\theta}u_{\infty}(t_{0})|_{H(t_{0})}^{2}\frac{\omega^{n}}{n!}\leq\check{C}.$ By (3.32), we know that $\hat{H}$ and $H(t_{0})$ are locally mutually bounded each other. By choosing a subsequence, we have $)u_{\infty}(t_{0})\rightarrow u_{\infty}\in L_{1}^{2}(\Gamma(S_{\hat{H}}$)) weakly in local $L_{1}^{2}$ on $X\setminus\Sigma$ as $t_{0}\rightarrow 0$, where $u_{\infty}$ satisfies
@@
(3.58) $\lim_{t_{0}\rightarrow 0}\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}(t_{0})\sqrt{-1}\Lambda_{\omega}F_{H(t_{0}),\theta_{X_{\mathrm{reg}}}})\frac{\omega^{n}}{n!}=\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta_{X_{\mathrm{reg}}}})\frac{\omega^{n}}{n!}.$
Set $\hat{u}_{\infty}(t_{0})=(h(t_{0}))^{\frac{1}{2}}\cdot u_{\infty}(t_{0})\cdot(h(t_{0}))^{-\frac{1}{2}}.$ It is easy to check that: $\hat{u}_{\infty}(t_{0})\in S_{\hat{H}}(\mathcal{E}|_{X\setminus\Sigma})$ and $|\hat{u}_{\infty}(t_{0})|_{\hat{H}}=|u_{\infty}(t_{0})|_{H(t_{0})}$. Following the same argument in [73, Lemma 4.2], we have:
@@
for any compact domain $\Omega\subset X\setminus\Sigma$ and any positive smooth function $\tilde{\Upsilon}:R\times R\rightarrow R$.
@@
By (3.56), (3.58) and (3.59), for any given compact domain $\Omega\subset X\setminus\Sigma$ and any given positive number $\tilde{\epsilon}>0$, we have
@@
(3.60) $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})\frac{\omega^{n}}{n!}+\int_{\Omega}\langle\Upsilon(\hat{u}_{\infty}(t_{0}))(\overline{\partial}_{\theta}\hat{u}_{\infty}(t_{0})),\overline{\partial}_{\theta}\hat{u}_{\infty}(t_{0})\rangle_{\hat{H}}\frac{\omega^{n}}{n!}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}+\tilde{\epsilon}$
@@
for small $t_{0}$. As we know that $\hat{u}_{\infty}(t_{0})\rightarrow u_{\infty}$ in $L^{2}(\Omega)$ as $t_{0}\rightarrow 0$, $|\hat{u}_{\infty}(t_{0})|_{\hat{H}}$ is uniformly bounded in $L^{\infty}$ and $|\overline{\partial}_{\phi}\hat{u}_{\infty}(t_{0})|_{\hat{H},\omega}$ is uniformly bounded in $L^{2}(\Omega)$. By the same argument as that in Simpson’s paper (Lemma 5.4 in [97]), we have
@@
(3.61) $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})\frac{\omega^{n}}{n!}+\|\Upsilon^{\frac{1}{2}}(u_{\infty})(\overline{\partial}_{\theta}u_{\infty})\|_{L^{q}(\Omega)}^{2}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}+2\tilde{\epsilon}$
for any $q<2$ and any $\tilde{\epsilon}>0$. Since $\tilde{\epsilon}$, $q<2$ and $\Omega$ are arbitrary, we get
@@
(3.62) $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})+\langle\Upsilon(u_{\infty})(\overline{\partial}_{\theta}u_{\infty}),\overline{\partial}_{\theta}u_{\infty}\rangle_{\hat{H}}\frac{\omega^{n}}{n!}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.$
@@
By the above inequality and the Lemma 5.5 in [97], we can see that the eigenvalues of $u_{\infty}$ are constant almost everywhere. Let $\mu_{1}<\dots<\mu_{l}$ denote the distinct eigenvalue of $u_{\infty}$. Since $\int_{X\setminus\Sigma}\operatorname{tr}u_{\infty}\frac{\omega^{n}}{n!}=0$ and $\|u_{\infty}\|_{L^{1}}=1$, we must have $l\geq 2$. By Uhlenbeck and Yau’s regularity theorem of $L_{1}^{2}$-subbundle ([104]) and Simpson’s result (page 887 in [97]), $\pi_{\alpha}$ represents a saturated $\theta_{\mathrm{reg}}$-invarient coherent sub-sheaf $\mathcal{E}_{X_{\mathrm{reg}},\alpha}$ of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ on the open set $X\setminus\Sigma$. Since $\mathcal{E}_{X}$ is reflexive, the trivial extension of $\mathcal{E}_{X_{\mathrm{reg}},\alpha}$ is a saturated subsheaf of $\mathcal{E}_{X}$ that $\theta_{X_{\mathrm{reg}}}$-invariant on $X_{\mathrm{reg}}$ by Lemma 2.5, which will be denoted by $\mathcal{E}_{\alpha}$. Let $\mathcal{E}_{\alpha,\mathrm{orb}}$ be the saturation of $\{\operatorname{image}(f_{\mu}^{*}\mathcal{E}_{\alpha}\xrightarrow{f_{\mu}^{*}}f_{\mu}^{*}\mathcal{E}_{X})\cap E_{\mu}\}$ on $E_{\mathrm{orb}}$. Recall Proposition 3.4, $f_{\mu}$ is étale outside $f_{\mu}^{-1}(\Sigma)$ and thus there exists an analytic subspace $Z\subset X$ of codimension $\geq 2$ such that $\mathcal{E}_{\alpha,\mathrm{orb}}:=\{f_{\mu}^{*}\mathcal{E}_{\alpha}/\mathrm{torsion}\}$. It follows from Lemma 2.18, Lemma 2.17 and the Chern-Weil formula in the context of orbifolds (cf. [110, Section 2.2.2]) that
$\deg_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{\alpha})=\deg_{f_{\mathrm{orb}}^{*}\omega}(\mathcal{E}_{\alpha,\mathrm{orb}})$

### v2

@@ 3.3.1. Donaldson’s functional
(3.38) $\|s\|_{P}=\sup_{X\setminus\Sigma}|s|+\|\bar{\partial}_{\theta}s\|_{L^{2}(X\setminus\Sigma)}+\|\Delta_{\omega,K}^{\prime}s\|_{L^{1}(X\setminus\Sigma)}.$
Let $\mathcal{P}$ denote the space of smooth Hermitian metrics $K$ such that $\int_{X\setminus\Sigma}|\Lambda_{\omega}F_{K,\theta}|_{\omega}\frac{\omega^{n}}{n!}<\infty.$ Due to Proposition 4.1 in [100], we can endow $\mathcal{P}$ with the structure of an analytic manifold with local charts $s\mapsto Ke^{s}$, where $s\in\mathcal{P}(S_{K})$. Let $\mathcal{P}_{0}$ be one of the components covered by these charts. Let us recall Donaldson’s functional defined on the space $\mathcal{P}_{0}$ (see Section 5 in [100] for details),
@@
The following proposition is standard because the argument of [100, Section 5] can be directly applied when $d\omega^{n-1}=0$.
@@
Let $H(t)=Kh(t)$ for $0\leq t\leq 1$ be a piecewise differentiable curve connecting metrics $K$ and $H$ in the space $\mathcal{P}_{0}$. If $d\omega^{n-1}\equiv 0$, then
(3.40) $\mu_{\omega}(K,H)=\int_{0}^{1}\int_{X\setminus\Sigma}\sqrt{-1}\operatorname{tr}(h^{-1}\frac{\partial h}{\partial t}\cdot F_{H(t),\theta})\wedge\frac{\omega^{n-1}}{(n-1)!}dt,$
@@
(3.41) $\frac{d}{dt}\mu_{\omega}(K,H(t))=\int_{X\setminus\Sigma}\sqrt{-1}\operatorname{tr}(h^{-1}\frac{\partial h}{\partial t}\cdot F_{H(t),\theta})\wedge\frac{\omega^{n-1}}{(n-1)!}$
@@ 3.3.2. Approximate Hermitian-Einstein structure
Let $H(t)$ be the long time solution of (3.9) on the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})|_{X\setminus\Sigma}$. Since we know that $|\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}|_{H(t)}$ is uniformly bounded for $t\geq t_{0}>0$, it follows from (3.37) and Lemma 3.18 that $H(t)$ (for every $t>0$) belongs to the space $\mathcal{P}_{0}$. According to (3.41), a formula for the derivative of Donaldson’s functional with respect to $t$ is given by
@@
(3.43) $r^{-\frac{1}{2}}\|s(t_{1},t_{2})\|_{L^{1}(X\setminus\Sigma,\omega,H(t_{1}))}\leq 2\tilde{C}_{1}(t_{2}-t_{1})+\operatorname{Vol}(X,\omega)\log(2r),$
(3.44) $\|s(t_{1},t_{2})\|_{L^{\infty}(X\setminus\Sigma,H(t_{1}))}\leq r^{\frac{1}{2}}\{2C_{S}\tilde{C}_{1}(t_{0}^{-n}+1)(t_{2}-t_{1})+\log 2r\}$
@@
(3.45) $\|s(t_{1},t_{2})\|_{L^{\infty}(X\setminus\Sigma,H(t_{1}))}\leq C_{1}(t_{0}^{-1})\|s(t_{1},t_{2})\|_{L^{1}(X\setminus\Sigma,\omega,H(t_{1}))}+C_{2}(t_{0}^{-1})$
for $0<t_{0}\leq t_{1}\leq t_{2}$, where $C_{1}(t_{0}^{-1})$ and $C_{2}(t_{0}^{-1})$ are constants depending only on $r$, the Sobolev constant $C_{S}$, $\tilde{C}_{1}$, $t_{0}^{-1}$ and $\operatorname{Vol}(X,\omega)$.
@@
(3.46) $\int_{X\setminus\Sigma}|\sqrt{-1}\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}-\lambda\operatorname{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}|_{H(t)}^{2}\frac{\omega^{n}}{n!}\leq\hat{C}_{1}^{\prime},$
where $\hat{C}_{1}^{\prime}$ is a positive constant. Define the limit:
(3.47) $\lim\limits_{t\rightarrow 0}\int_{X\setminus\Sigma}|\sqrt{-1}\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}-\lambda\operatorname{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}|_{H(t)}^{2}\frac{\omega^{n}}{n!}=C^{\ast}\geq 0.$
Assume further that there exists a sequence $t_{i}\rightarrow\infty$ such that
@@
Then the following statement holds: there is a section $)u_{\infty}\in L_{1}^{2}(\Gamma(S_{\hat{H}}$)) such that $\operatorname{tr}u_{\infty}=0$, $\|u_{\infty}\|_{L^{2}}=1$ and the eigenvalues of $u_{\infty}$ are constants almost everywhere and not all equal. Furthermore, assume $\nu_{1}<\nu_{2}<\cdots<\nu_{l}$ ($l\geq 2$) are the distinct eigenvalues of $u_{\infty}$. For each $1\leq k\leq l-1$, define a smooth function $P_{k}:\mathbb{R}\to\mathbb{R}$ such that
$P_{k}(x)=\left\{\begin{aligned} \ 1,\quad&x\leq\mu_{k},\\ \ 0,\quad&x\geq\mu_{k+1}.\end{aligned}\right$.
Set $\pi_{k}=P_{k}(u_{\infty})$. Then each $\pi_{k}$ defines a saturated $\theta_{X_{\mathrm{reg}}}$-invariant subsheaf $\mathcal{E}_{X_{\mathrm{reg}},k}$ of the Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$, and each $\mathcal{E}_{X_{\mathrm{reg}},k}$ extends to a saturated subsheaf $\mathcal{E}_{k}$ of $\mathcal{E}_{X}$ such that
(3.49) $2\pi\sum_{k=1}^{l-1}(\mu_{k+1}-\mu_{k})\operatorname{rank}(\mathcal{E}_{k})(\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})-\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{k}))\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}^{\prime}}.$
@@
(3.50) $\|s(t_{0},t_{i})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}\rightarrow+\infty,$
@@
for all $0<t_{0}\leq 1\leq t$, where $C_{3}$ and $C_{4}$ are constants depending only on the Sobolev constant $C_{S}$, $r$, $\hat{C}_{1}^{\prime}$ and the geometry of $(X\setminus\Sigma,\omega)$.
Define $u_{i}(t_{0})=\|s(t_{0},t_{i})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}^{-1}s(t_{0},t_{i})\in S_{H(t_{0})}$, then $\|u_{i}(t_{0})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}=1$ and
@@
(3.54) $\liminf_{t\rightarrow+\infty}\frac{-\mu_{\omega}(H(t_{0}),H(t))}{\|s(t_{0},t)\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}}\geq r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}^{\prime}}.$
From the inequalities (3.54), (3.48), (3.51) and Lemma 5.4 in [100], it follows that, after passing to a subsequence still denoted by $\{u_{i}(t_{0})\}$, we have $u_{i}(t_{0})\rightharpoonup u_{\infty}(t_{0})$ weakly in $L_{1}^{2}$ as $i\rightarrow\infty$, where the limit $u_{\infty}(t_{0})$ satisfies: $\|u_{\infty}(t_{0})\|_{L^{1}(X\setminus\Sigma,\omega,H(t_{0}))}=1$, $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}(t_{0}))\frac{\omega^{n}}{n!}=0$ and
@@
Furthermore, if $\Upsilon:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ is a positive smooth function such that $\Upsilon(\lambda_{1},\lambda_{2})<(\lambda_{1}-\lambda_{2})^{-1}$ whenever $\lambda_{1}>\lambda_{2}$, then
(3.56) $\begin{split}&\int_{X\setminus\Sigma}\big(\operatorname{tr}(u_{\infty}(t_{0})\sqrt{-1}\Lambda_{\omega}(F_{H(t_{0}),\theta}))+\langle\Upsilon(u_{\infty}(t_{0}))(\overline{\partial}_{\theta}u_{\infty}(t_{0})),\overline{\partial}_{\theta}u_{\infty}(t_{0})\rangle_{H(t_{0})}\big)\frac{\omega^{n}}{n!}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}^{\prime}}.\end{split}$
Clearly (3.14) and (3.55) mean that $\|\Lambda_{\omega}(F_{H(t_{0}),\theta})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}$ and $\|u_{\infty}(t_{0})\|_{L^{\infty}(X\setminus\Sigma,H(t_{0}))}$ are uniformly bounded (independent of $t_{0}$). Therefore, (3.56) yields that there exists a uniform constant $\check{C}^{\prime}$ independent of $t_{0}$ such that $\int_{X\setminus\Sigma}|\overline{\partial}_{\theta}u_{\infty}(t_{0})|_{H(t_{0})}^{2}\frac{\omega^{n}}{n!}\leq\check{C}^{\prime}.$ From (3.32), we know that $\hat{H}$ and $H(t_{0})$ are locally mutually bounded. By choosing a subsequence, we obtain $)u_{\infty}(t_{0})\rightarrow u_{\infty}\in L_{1}^{2}(\Gamma(S_{\hat{H}}$)) weakly in local $L_{1}^{2}$ on $X\setminus\Sigma$ as $t_{0}\rightarrow 0$, where $u_{\infty}$ satisfies
@@
(3.58) $\lim_{t_{0}\rightarrow 0}\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}(t_{0})\sqrt{-1}\Lambda_{\omega}F_{H(t_{0}),\theta})\frac{\omega^{n}}{n!}=\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})\frac{\omega^{n}}{n!}.$
Set $\hat{u}_{\infty}(t_{0})=(h(t_{0}))^{\frac{1}{2}}\cdot u_{\infty}(t_{0})\cdot(h(t_{0}))^{-\frac{1}{2}}.$ One can verify that: $\hat{u}_{\infty}(t_{0})\in S_{\hat{H}}(\mathcal{E}|_{X\setminus\Sigma})$ and $|\hat{u}_{\infty}(t_{0})|_{\hat{H}}=|u_{\infty}(t_{0})|_{H(t_{0})}$. Following the same argument as in the proof of [74, Lemma 4.2], we deduce
@@
for any compact domain $\Omega\subset X\setminus\Sigma$ and any positive smooth function $\tilde{\Upsilon}$$\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$.
@@
Combining (3.56), (3.58) with (3.59), for any given compact domain $\Omega\subset X\setminus\Sigma$ and any $\tilde{\epsilon}>0$, we get
@@
(3.60) $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})\frac{\omega^{n}}{n!}+\int_{\Omega}\langle\Upsilon(\hat{u}_{\infty}(t_{0}))(\overline{\partial}_{\theta}\hat{u}_{\infty}(t_{0})),\overline{\partial}_{\theta}\hat{u}_{\infty}(t_{0})\rangle_{\hat{H}}\frac{\omega^{n}}{n!}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}^{\prime}}+\tilde{\epsilon}$
@@
for sufficiently small $t_{0}$. Since $\hat{u}_{\infty}(t_{0})\rightarrow u_{\infty}$ in $L^{2}(\Omega)$ as $t_{0}\rightarrow 0$, with $|\widehat{u}_{\infty}(t_{0})|_{\widehat{H}}$ uniformly bounded in $L^{\infty}$ and $|\bar{\partial}_{\theta}\widehat{u}_{\infty}(t_{0})|_{\widehat{H},\omega}$ uniformly bounded in $L^{2}(\Omega)$, the same argument as in [100, Lemma 5.4] yields
@@
(3.61) $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})\frac{\omega^{n}}{n!}+\|\Upsilon^{\frac{1}{2}}(u_{\infty})(\overline{\partial}_{\theta}u_{\infty})\|_{L^{q}(\Omega)}^{2}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}^{\prime}}+2\tilde{\epsilon}$
for any $q<2$. Since $\tilde{\epsilon}$, $q<2$ and $\Omega$ are arbitrary, we get
@@
(3.62) $\int_{X\setminus\Sigma}\bigg(\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})+\langle\Upsilon(u_{\infty})(\overline{\partial}_{\theta}u_{\infty}),\overline{\partial}_{\theta}u_{\infty}\rangle_{\hat{H}}\bigg)\frac{\omega^{n}}{n!}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}^{\prime}}.$
@@
From the above inequality and [100, Lemma 5.5], we can see that the eigenvalues of $u_{\infty}$ are constant almost everywhere. Let $\nu_{1}<\dots<\nu_{l}$ denote the distinct eigenvalues of $u_{\infty}$. Since $\int_{X\setminus\Sigma}\operatorname{tr}u_{\infty}\frac{\omega^{n}}{n!}=0$ and $\|u_{\infty}\|_{L^{1}}=1$, we must have $l\geq 2$. By Uhlenbeck and Yau’s regularity theorem for $L_{1}^{2}$-subbundle [107] and Simpson’s result [100, Page 887], each $\pi_{k}$ represents a saturated $\theta_{\mathrm{reg}}$-invarient coherent subsheaf $\mathcal{E}_{X_{\mathrm{reg}},k}$ of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ on the open set $X\setminus\Sigma$. Since $\mathcal{E}_{X}$ is reflexive, the trivial extension of $\mathcal{E}_{X_{\mathrm{reg}},k}$ is a saturated subsheaf of $\mathcal{E}_{X}$ that is $\theta_{X_{\mathrm{reg}}}$-invariant on $X_{\mathrm{reg}}$ by Lemma 2.5, which will be denoted by $\mathcal{E}_{k}$. Let $\mathcal{E}_{k,\mathrm{orb}}$ be the saturation of $\{\operatorname{image}(f_{\alpha}^{*}\mathcal{E}_{k}\xrightarrow{f_{\alpha}^{*}}f_{\alpha}^{*}\mathcal{E}_{X})\cap E_{\alpha}\}$ in $E_{\mathrm{orb}}$. Recall from Proposition 3.4 that $f_{\alpha}$ is étale outside $f_{\alpha}^{-1}(\Sigma)$ and thus there exists an analytic subspace $Z\subset X$ of codimension $\geq 2$ such that $\mathcal{E}_{k,\mathrm{orb}}=\{f_{\alpha}^{*}\mathcal{E}_{k}/(\mathrm{torsion})\}$ on $Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(Z)$. It follows from Lemma 2.18, Lemma 2.17 and the Chern-Weil formula for orbifolds (cf. [113, Section 2.2.2]) that
$\deg_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{k})=\deg_{f_{\mathrm{orb}}^{*}\omega}(\mathcal{E}_{k,\mathrm{orb}})$